OpenStudy (anonymous):

/16x2y8 (x to the second power y to the eight power)?

OpenStudy (anonymous):

What's your question, though?

OpenStudy (anonymous):

@lokistan divide 16x2y8 (x to the second power y to the eight power)?

OpenStudy (anonymous):

Divide it by what, though?

OpenStudy (anonymous):

Are you dividing 16x^2 BY y^8 or something?

OpenStudy (anonymous):

a pice of wire 5 /3 m long is to be cut into 15 pieces of the same length. what is the length of each piece

OpenStudy (anonymous):

\[\frac{\frac{5}{3}}{15}=\frac{5}{3}\times \frac{1}{15}=\frac{5}{3}\times \frac{1}{3 \times 5}=\frac{1}{3 \times 3} = \frac{1}{9}\]

OpenStudy (anonymous):

e23, what exactly does your question say?

OpenStudy (anonymous):

@ lokisan it doesnt tell me anything it just has the problem...im confused :/ and I have to choose between a) 4xy2 (y to the second power) b) 4xy4 (y to the fourth power) c)8xy4 (y to the fourth power) d)8xy6 (y to the sixth power) as an answer

OpenStudy (anonymous):

I'm confused too, because I can't help you choose an answer without a problem :(

OpenStudy (anonymous):

You're given a choice of several answers, so are you saying there is no question beforehand?

OpenStudy (anonymous):

yes it just has the problem then a choice of answers but I don't know what I'm dividing....thanks for your help, I'll ask my teacher to explain

OpenStudy (anonymous):

OK. Good luck.